OpenStudy (samsan9):

the foci for the hyperbola (x-2)^2/36-(y+1)^2/64=1 is (2, -1+4sqrt(7) and (2,-1-4sqrt(7) true or false

OpenStudy (samsan9):

how do you know what the foci is?

OpenStudy (samsan9):

\[(2,-1+4\sqrt{7}) and (2,-1-4\sqrt{7})\]

OpenStudy (mathmate):

\(\large \frac{x^2}{a^2}-\frac{y^2}{b^2}=1\) hyperbola with axis on x-axis, centred at origin

OpenStudy (mathmate):

Foci are at \(\sqrt{a^2+b^2}\) from the origin.

OpenStudy (mathmate):

Do you know where the centre of this hyperbola is?

OpenStudy (samsan9):

(2,-1) i think

OpenStudy (mathmate):

Good! Does it have a horizontal or vertical axis?

OpenStudy (samsan9):

idk :/ but i think its horizontal

OpenStudy (mathmate):

Yes, because \(\large \frac{x^2}{a^2}-\frac{y^2}{b^2}=-1\) hyperbola with axis on y-axis, centred at origin

OpenStudy (mathmate):

So what are a, b and c=\(\sqrt{a^2+b^2}\)?

OpenStudy (samsan9):

c= a=2 and b=-1

OpenStudy (samsan9):

idk tbh

OpenStudy (mathmate):

If you compare the equation of the standard ellipse with horizontal axis, you will find a^2=36, b^2=64, giving a=6, b=8, c=10. And since the axis is horizontal, and hyperbola is centred at (h,k)=(2,-1) we have the foci (h+c,k), (h-c,k)